OpenStudy (anonymous):

Use the rule y = -x + 7 to fill in the blank. If (-1, y) is on the graph, then y =

OpenStudy (anonymous):

Oh... actually its 8... this one is -1 not 1.

OpenStudy (anonymous):

Thank you! You've helped alot

OpenStudy (anonymous):

np :)

OpenStudy (gorv):

put x=-1 y=-(-1)+7=1+7=8

OpenStudy (anonymous):

Use the function f(x) = 2 - 5x to fill in the blanks in the table. x -2 -1 0 1 2

OpenStudy (anonymous):

And thx

OpenStudy (anonymous):

2-5(-2) = 2+10 = 12 2-5(-1) = 2+5 = 7 2-5(0) = 2-0 = 2 2-5(1) = 2-5 = -3 2-5(2) = 2-10 = -8

OpenStudy (anonymous):

Sorry computer's very slow

OpenStudy (anonymous):

Thanks